OpenStudy (anonymous):

integrate x^2 dx / (xsinx+cosx )^2

OpenStudy (helder_edwin):

\[ \large \int\frac{x^2}{x\sin x+\cos x}\,dx \] ??

OpenStudy (turingtest):

\[\large \int\frac{x^2}{(x\sin x+\cos x)^2}\,dx\]I think...

OpenStudy (turingtest):

even worse

OpenStudy (helder_edwin):

right i forgot the squared

OpenStudy (turingtest):

not an ugly answer, but no steps from wolfram

hartnn (hartnn):

xsin x +cos x is derivative of x sin x will this help?

OpenStudy (turingtest):

\[x\cos x + \sin x=\frac d{dx}(x\sin x)\]

OpenStudy (turingtest):

so no, I guess not